OpenStudy (anonymous):

-5x + y = -2 -3x + 6y = -12

hartnn (hartnn):

substitution again ?

OpenStudy (anonymous):

Yes ma'am.

OpenStudy (skullpatrol):

You closed the last question before you checked it?

hartnn (hartnn):

add 5x on both sides of 1st equation, what you get ? its pretty much same as previous.

OpenStudy (anonymous):

@skullpatrol what? Im new here. d:

hartnn (hartnn):

i think you try this first on your own, and if you don't get then ask...

OpenStudy (anonymous):

@hartnn is it going to y= 5x-2?

OpenStudy (anonymous):

x is pretty clearly 0 !

hartnn (hartnn):

now put y=5x-2 in 2nd equation

OpenStudy (anonymous):

Is it going to be 30x?

OpenStudy (anonymous):

sorry guys, whats the question is?

OpenStudy (anonymous):

find x and y yeah?

hartnn (hartnn):

-3x + 6y = -12 -3x+6(5x-2)=-12 tell me what you get as entire eqution, not just 30x.

OpenStudy (anonymous):

-3x + 30x + 12 = -12

hartnn (hartnn):

6*(-2) =... ?

hartnn (hartnn):

+12 or -12 ??

OpenStudy (anonymous):

find the corect answer with substition aye4angela

OpenStudy (anonymous):

@hartnn -12

hartnn (hartnn):

then you get -3x + 30x -12 = -12 right ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

now add 12 on both sides, what you get ?

OpenStudy (anonymous):

0

hartnn (hartnn):

good! x=0 is correct :) so, y= .. ?

OpenStudy (anonymous):

-2?

hartnn (hartnn):

correct again! now you are good at this :)

OpenStudy (anonymous):

Ohhhhhh.(: i feel smart. ^-^ well that was the last problem. :3 thank you so very much!(:

hartnn (hartnn):

welcome very much! ^_^